# CORE: Lossless Compression for Retrieval-Augmented LLMs via Reinforcement Learning

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the timeliness of knowledge and the factual accuracy of responses in Large Language Models (LLMs). However, the inclusion of excessive retrieved documents substantially increases the input length, leading to higher computational costs. Previous studies have attempted to compress retrieved documents into shorter texts before in-context integration, but such methods often compromise end-task performance. The lack of well-defined compression targets forces many approaches to rely on fixed heuristics, which cannot guarantee that the compressed content will effectively support the end task. To address these limitations, we propose CORE, a novel method designed to achieve lossless context compression for RAG. CORE employs reinforcement learning to optimize the compression process without relying on predefined compression labels. Specifically, it utilizes end-task performance as a reward signal and applies Generalized Reinforcement Learning Policy Optimization (GRPO) to train the compressor. This end-to-end training framework enables the compressor to generate summaries that maximize the accuracy of answers generated by the LLM. Extensive experiments on four datasets demonstrate the superiority of our approach. With a high compression ratio of 3%, our method not only avoids performance degradation compared to prepending full documents across all datasets but also improves the average Exact Match (EM) score by 3.3 points. The code will be released soon.
@@ 1 Introduction
Large language models (LLMs) have undergone rapid development in recent years, significantly enhancing performance across various language tasks due to their emergent capabilities in semantic understanding and reasoning. Nevertheless, LLMs still face challenges in updating knowledge and providing factual responses Fan et al. (2024). To address these issues, Retrieval-Augmented Generation (RAG) has emerged as a promising approach Ram et al. (2023). By retrieving the most relevant documents from extensive knowledge bases and prepending them as contextual information to the original input, RAG substantially improves LLM performance on knowledge-intensive tasks Ram et al. (2023); Fan et al. (2024).
As shown in Figure 1, in-context retrieval-augmented generation significantly outperforms the baseline without RAG (i.e., the number of docs equals to zero). As the number of prepended retrieved documents increases, the accuracy continues to rise, eventually saturating when the number of documents reaches 10. However, this improvement in accuracy introduces two limitations: a substantial rise in computational costs due to the increased number of tokens that must be encoded, and challenges for the model in effectively utilizing all information in the context, often resulting in the omission of key information located in the middle Liu et al. (2023).
To mitigate these issues, recent research has made strides in compressing retrieved documents Xu et al. ; Cao et al. (2024); Jin et al. (2024b); Zhu et al. (2024); Kim & Thorne (2025); Wu et al. (2025); Jin et al. (2024a); Zhang et al. (2024a). These methods mainly include summarizing retrieved documents Xu et al. , extracting key information Xu et al. ; Cao et al. (2024), constructing key supporting evidence Jin et al. (2024b), and filtering noise based on information theory Zhu et al. (2024). Although these efforts have shown promising progress, they still face several limitations. Firstly, most methods achieve document compression at the expense of performance degradation; for instance, RECOMP Xu et al. exhibits a decline of 3–5 exact match (EM) points after compression (Table 1), making it unsuitable for accuracy-sensitive scenarios. Secondly, despite serving downstream tasks, the absence of ideal compression targets prevents most existing methods from being optimized in an end-to-end manner. This limitation hinders the compressor’s ability to generate content that reliably supports the LLM in producing accurate responses, thereby undermining end-task performance. Finally, some approaches employ compressors with parameter counts comparable to those of the target LLM itself, resulting in non-negligible computational costs that may offset the efficiency gains of compression.
To overcome these limitations, we propose COmpression via REinforcement learning (CORE), a novel method designed to achieve lossless context compression for retrieval-augmented generation. Recently, reinforcement learning (RL) has emerged as a promising approach for enhancing reasoning capabilities without requiring supervised data on intermediate reasoning steps Liu et al. (2024); Chen et al. (2025a). Inspired by this advancement, we employ RL to train a compressor that has significantly fewer parameters than the main LLM to summarize documents. Specifically, the summaries generated by the compressor are concatenated with the original input and fed into the LLM to obtain predicted answers. A rule based reward signal is then formulated based on answer accuracy, i.e., end task performance, to optimize the compressor. The training process is implemented using the Group Policy Optimization (GRPO) Shao et al. (2024) algorithm, which effectively integrates the reward signal derived from end-task performance to update the compressor in a stable and efficient manner. This end-to-end design enables the compressor to be trained in a goal-oriented manner even in the absence of ideal compression labels, ensuring that the generated summaries effectively assist the LLM in producing accurate responses.
@@
We evaluate the effectiveness of CORE on four benchmark datasets: two single-hop question-answering datasets, Natural Questions and TriviaQA, along with two multi-hop datasets, HotpotQA and 2WikiMultihopQA. Experimental results (Table 1) demonstrate that CORE achieves state-of-the-art performance compared to all baselines. While maintaining a high compression ratio of 3%, our method not only avoids performance degradation relative to prepending full documents across all four datasets, but also improves the average Exact Match (EM) score by 3.3 points. Furthermore, we empirically show that our approach is not dependent on a specific compressor architecture (Figure 5). The trained compressor also exhibits strong transferability to other large language models (Table 2). Finally, via a detailed case study (Table 4 and 5), we highlight the advantages of our compression method over alternative approaches.
Figure 1: Growth of EM and F1 scores with increasing retrieved documents on two datasets.
@@ 2 CORE-RAG
Figure 2: Overview of our method.
@@
In this section, we introduce our proposed method, CORE-RAG. First, we provide an overview of the entire workflow. Then, we detail how to train a compressor using reinforcement learning to significantly reduce the number of document tokens while maintaining performance.
2.1 Pipeline
@@
We adopt the same problem formulation with prior work Xu et al. , where given an input sequence $x$, target output sequence $y$, and a set of $N$ retrieved documents $D$, our objective is to compress $D$ with respect to $x$ into a summary $s$ that preserves the most relevant information while using significantly fewer tokens than $D$. The whole process consists of two key components: a compression model $c_{\theta}\colon(x,D)\mapsto s$ and a large language model $M\colon(x,s)\mapsto y$ that generates the output sequence $y$. We treat $M$ as a black-box system while focusing exclusively on training the compressor. The compressor employed in this study is also a language model. However, it is intentionally designed to be significantly smaller than $M$, as we aim to reduce computational costs of encoding the retrieved documents.
2.2 Training the Compressor with RL
Our compressor is designed to generate document summaries while ensuring that these summaries are particularly useful for the base LLM to perform a downstream task. The latter requirement is extremely challenging as the criteria for an effective summary is inherently task-dependent, meaning that no supervised training data is available to guide the compression. To address this challenge, we employ reinforcement learning, motivated by its demonstrated success in training language models without explicit intermediate supervision. The overview of our training method is shown in Figure 2. The remainder of this section elaborates on our RL training methodology, including distillation warm-up, policy optimization, and reward formulation.
2.2.1 Distillation for RL Warm-Start
@@
Since our chosen compression model has a relatively small number of parameters, its query-focused document summarization capability is limited. To provide reinforcement learning with a strong initial policy and ensure training stability, we first employ knowledge distillation from a teacher model to initialize the RL policy. Specifically, we use an extremely large language model (DeepSeek-V3) as the teacher model to generate summaries of the retrieved document sets in response to the given query, which serve as target summaries. We then evaluate the end task performance under two conditions: (1) with the target summary prepended to the input, and (2) with the original input alone (i.e., no retrieval). By comparing these results on the base model ($M$), we filter out instances where prepending the summary does not improve performance. The remaining samples are used for supervised fine-tuning.
2.2.2 Group Relative Policy Optimization
After the warm-up phase, we employ reinforcement learning to train the compressor in an end-to-end manner. Our objective is to optimize the summaries generated by the compressor to improve the performance of the end task. However, obtaining perfect summary labels is challenging, as even summaries generated by extremely large language models cannot be guaranteed to be the most effective. Therefore, we leverage reinforcement learning to optimize the compressor, allowing it to learn how to generate summaries that enhance end-task performance.
@@ 2.2.2 Group Relative Policy Optimization
Specifically, we employ GRPO Shao et al. (2024) as the learning algorithm. Unlike Proximal Policy Optimization (PPO), which trains a separate critic model, GRPO estimates the baseline directly from a group of rollouts. Given an existing policy, $\pi_{\theta_{\text{old}}}$, and a reference policy, $\pi_{\theta_{\text{ref}}}$, the GRPO objective maximizes the policy $\pi_{\theta}$ using $G$ rollouts $\tau=\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)$, for each input $x\sim\mathcal{D}$:
$\mathcal{J}(\theta)$ $={\mathbb{E}}_{x\sim\mathcal{D},\ \{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}$ (1)
@@
$\frac{1}{G}\sum_{i=1}^{G}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\ \text{clip}\Bigg{(}\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\epsilon,1+\epsilon\Bigg{)}A_{i}\Bigg{)}-\beta\mathbb{D}_{\text{KL}}\big{(}\pi_{\theta}\parallel\pi_{\theta_{\text{ref}}}\big{)}\Bigg{]},$
@@
2.2.3 Reward
  • Generating End-Task Output. It is important to note that the reward is not calculated directly based on the output summary of the compressor. Instead, the summary generated by the compressor is prepended to the original input (question) and then fed into the base LLM $M$ to produce an output $y$. The reward is subsequently calculated based on $y$. Throughout the entire training process, $M$ is fixed and does not participate in training.
@@ 2.2.3 Reward
    • EM Reward ($r_{\text{EM}}$). We employ exact match (EM) as a main reward function, which is a widely adopted metric for evaluating the accuracy of generated outputs. The EM reward is binary: it yields a value of 1 if the generated answer perfectly matches the ground truth, and 0 otherwise.
    • F1 Reward ($r_{\text{F1}}$). Since exact matches occur infrequently in practice, relying solely on EM-based rewards leads to sparse reward signals. Furthermore, the EM metric fails to distinguish between partially correct answers, as all non-exact matches receive zero reward. Therefore, we introduce F1 score, which provides finer-grained evaluation by measuring the degree of partial match.
    The final reward function combines these signals through weighted summation:
@@
    where $\alpha\in(0,1]$ is a hyperparameter that controls the relative contribution of the F1 metric.
@@ 3.1 Experimental Settings
  • Datasets. We evaluate our method on four benchmark datasets: two single-hop question-answering datasets, Natural Questions Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017), and two multi-hop question-answering datasets, HotpotQA Yang et al. (2018) and 2WikiMultihopQA Ho et al. (2020). Results are reported on the test sets of Natural Questions and TriviaQA, as well as the development sets of HotpotQA and 2WikiMultihopQA. Performance is measured using Exact Match and token-level F1 scores for answer strings.
  • Base Large Language Models. We use Qwen2.5-14B-Instruct as the primary LLM model $M$ to generate predicted answers which are used to guide the training of the compressor. To evaluate the generalization ability of our method, we also transfer to another LLM model, LLama3.1-8B-Instruct.
@@
  • Compressor Models. We trained our compressor model using Qwen2.5-1.5B-Instruct to generate summaries of the retrieved documents. To evaluate the effect of using different models for compression, we also trained compressors based on Llama3.2-1B-Instruct and Llama3.2-3B-Instruct.
@@
  • Retrieval Corpus and Retrievers. Following previous studies Xu et al. , we use the Wikipedia corpus from December 20, 2018, as the retrieval source for all four datasets. The articles are segmented into non-overlapping 100-word documents. To ensure that our method is not dependent on a specific retriever, we experiment with several mainstream retrieval models. Specifically, we use DPR Karpukhin et al. (2020) for the Natural Questions dataset, a hybrid of DPR and BM25 Robertson et al. (1995) for the TriviaQA dataset, and the Contriever model Izacard et al. (2021) trained on the MS MARCO dataset Nguyen et al. (2016) for HotpotQA and 2WikiMultihopQA.
    Figure 3: Prompt template used to generate the summary.
    Figure 4: Prompt template for LLM QA.
  • Prompt Format. The prompt for the compressor is shown in Figure 3, which is used to generate summary of documents. For the question-answering prompt provided to the LLM, we include few-shot in-context examples in the prompt, followed by the retrieved documents and the question, following Xu et al. . Three in-context examples are randomly sampled from the training set. To remain consistent with Xu et al. , the retrieved documents are concatenated in ascending order of their retrieval scores, with the highest-scoring document placed closest to the question Si et al. (2022). This prompt template is provided in Figure 4.
@@
  • Baselines. To evaluate the effectiveness of our method, we compared it against various baselines. First, we evaluated the uncompressed approach—retaining the original in-context RALM setup—by prepending the top 1, 3, 5, and 10 retrieved documents to the prompt. We also tested alternative compression methods, including the traditional BM25 algorithm (which ranks sentences by their similarity to the input), DeepSeek-V3 model (671B parameters, far exceeding our compressor’s capacity), and the off-the-shelf Qwen2.5-1.5B-Instruct model (with comparable parameter size to our approach). Additionally, we included state-of-the-art context compression methods for RAG, RECOMP Xu et al. and NoiseFilter-IB Zhu et al. (2024). For RECOMP, we evaluate both its abstractive (summary-based) and extractive (sentence selection) variants.
@@
  • Implementation Details. The reinforcement learning framework in our study is based on verl  [^1] https://github.com/volcengine/verl. We train the compression model for two epochs on each dataset. Training is conducted on eight NVIDIA H20 GPUs using full parameter optimization, with a learning rate of 1e-5, a batch size of 256, five rollouts, and a KL loss coefficient of 0.001. For the base LLM used to generate rewards, we deploy it using vLLM. For the distillation warm-up phase, we perform full-parameter supervised fine-tuning on the compression model for two epochs using LLaMA-Factory  [^2] https://github.com/hiyouga/LLaMA-Factory.
@@
Table 1: Open-domain QA results using Qwen2.5-14B-Instruct as the LLM. The reported token counts represent the length of in-context evidence documents, excluding few-shot examples. RECOMP, NoiseFilter-IB and our proposed CORE are all trained using Qwen2.5-1.5B-Instruct.

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the timeliness of knowledge updates and the factual accuracy of responses in large language models. However, incorporating a large number of retrieved documents significantly increases input length, leading to higher computational costs. Existing approaches to document compression tailored for RAG often degrade task performance, as they typically rely on predefined heuristics in the absence of clear compression guidelines. These heuristics fail to ensure that the compressed content effectively supports downstream tasks. To address these limitations, we propose CORE, a novel method for lossless context compression in RAG. CORE is optimized end-to-end and does not depend on predefined compression labels, which are often impractical to obtain. Instead, it leverages downstream task performance as a feedback signal, iteratively refining the compression policy to enhance task effectiveness. Extensive experiments across four datasets demonstrate the effectiveness of CORE. With a high compression ratio of 3%, CORE not only prevents performance degradation compared to including full documents (i.e., without compression) but also improves the average Exact Match (EM) score by 3.3 points. The code for CORE will be released soon.
@@ 1 Introduction
Large language models (LLMs) have undergone rapid development in recent years, significantly enhancing performance across various language tasks due to their emergent capabilities in semantic understanding and reasoning. Nevertheless, LLMs still face challenges in updating knowledge and providing factual responses (Fan et al., 2024). To address these issues, Retrieval-Augmented Generation (RAG) has emerged as a promising approach. By retrieving the most relevant documents from external knowledge bases and prepending them as contextual information to the original input, RAG substantially improves LLM performance on knowledge-intensive tasks (Ram et al., 2023).
While RAG enhances performance, its effectiveness is closely tied to the number of retrieved documents used, since a broader context increases the probability of encompassing critical evidence. As illustrated in Figure 1, performance was weakest without any retrieved documents (i.e., without RAG). Accuracy improved consistently as more documents were added to the context, ultimately exceeding the no-RAG baseline by over 10 Exact Match (EM) points. However, this performance gain came with two significant limitations: (1) a substantial increase in computational cost from processing a larger number of context tokens (Xu et al., 2024), and (2) the model’s difficulty in effectively leveraging all provided documents, often resulting in the omission of key information located in the middle of the context (Liu et al., 2023).
These limitations have motivated recent research efforts aimed at compressing the retrieved context (Jin et al., 2024b; Wu et al., 2025; Jin et al., 2024a; Zhang et al., 2024a). Prominent approaches include document summarization (Xu et al., 2024), key information extraction (Cao et al., 2024; Xu et al., 2024), the construction of key supporting evidence (Jin et al., 2024b), and noise filtering based on information theory (Zhu et al., 2024). Despite recent progress, these methods have several notable shortcomings. First, compression often results in a performance trade-off. For instance, RECOMP (Xu et al., 2024) suffers a 3–5 point drop in EM score (Table 1), making it unsuitable for accuracy-sensitive applications. Second, most compression methods are heuristic in nature. The models are typically trained to generate summaries that are generally good but not necessarily useful for the downstream answer-generation LLM. This limitation arises from the lack of an ideal supervisory signal that defines what an optimal summary should be for the downstream task. This fundamental gap hinders the end-to-end optimization of existing methods. Finally, some compression models (Zhu et al., 2024) have parameter counts comparable to the LLM that performs the end task, resulting in substantial computational costs that undermine the efficiency gains of compression.
Addressing these critical shortcomings requires a method that aligns compression with downstream task requirements, thereby minimizing performance trade-offs. To bridge this gap, we propose CORE, a novel method designed to achieve lossless context compression for RAG. Unlike previous compression methods, CORE is optimized in an end-to-end manner. Since obtaining predefined summary labels for supervision is impractical, we instead use downstream task performance as a feedback signal to evaluate the compression model’s output. This feedback enables iterative refinement of the compression policy, guiding the compression model toward improved downstream performance. To this end, we employ Group Relative Policy Optimization (GRPO), a technique particularly well-suited for this purpose (Liu et al., 2024; Shao et al., 2024; Chen et al., 2025a). In our framework, the accuracy of the downstream QA task is formalized as a reward, with the compression policy optimized through group-wise relative comparisons. Furthermore, our compression model is substantially smaller than the downstream LLM, which significantly reduces the computational overhead associated with encoding retrieved documents.
@@
We evaluate CORE on four benchmark datasets: two single-hop QA datasets (Natural Questions and TriviaQA) and two multi-hop datasets (HotpotQA and 2WikiMultihopQA). As shown in Table 1, CORE achieves state-of-the-art performance across all baselines. With a compression ratio of 3%, our approach not only avoids performance degradation compared to prepending full documents but also improves the average EM score by 3.3 points. We further demonstrate two key advantages of our approach. First, the effectiveness of CORE is not tied to a specific model architecture, as demonstrated by the fact that lossless compression can be achieved when various models are trained as compressors (Figure 3). Second, the compressor exhibits strong transferability: a compression model trained using feedback from one LLM generalizes effectively to other LLMs (Table 2). Finally, an in-depth case study (Tables 4 and 5) provides a qualitative analysis of the benefits of CORE.
Figure 1: Performance evolution with an increasing number of retrieved documents on two datasets. Traditional RAG requires more documents for better performance, while our method achieves comparable or superior results with significant token compression.
@@ 2 CORE-RAG
Figure 2: Overview of our method CORE. The upper section illustrates the inference pipeline. The lower section depicts the end-to-end training method for the compression model.
@@
This section introduces our proposed method, COmpression via REinforcement learning (CORE), which is shown in Figure 2. First, we provide an overview of the entire workflow. Then, we detail the end-to-end training strategy for the compression model, which is designed to drastically reduce the number of document tokens while preserving task performance.
2.1 Problem Formulation
@@
We adopt the same problem formulation as prior work (Xu et al., 2024). Given an input question $q$, a target output $y$, and a set of $k$ retrieved documents $D$, our objective is to compress $D$ with respect to $q$ into a summary $s$ that preserves the most useful information while using significantly fewer tokens than $D$. This summary $s$ is then prepended to the original input $q$ and fed into an LLM to generate the final answer for the downstream task. This overall pipeline is illustrated in the upper half of Figure 2. The process involves two key components: a compression model $\pi_{\theta}\colon(q,D)\mapsto s$ and a large language model $M\colon(s,q)\mapsto\hat{y}$, which generates the predicted answer $\hat{y}$. We treat $M$ as a black-box system and focus exclusively on training the compressor $\pi_{\theta}$. The compressor itself is also a language model, but it is intentionally designed to be significantly smaller than $M$ to reduce the computational cost of encoding the retrieved documents.
2.2 Training the Compressor
Our compressor is designed to generate document summaries that are highly useful to the LLM ($M$) for downstream tasks. This objective is challenging because the criteria for an effective summary are task-dependent, and direct supervision is unavailable. We therefore formulate this as an end-to-end training problem and employ reinforcement learning to optimize the compressor without relying on pre-defined compression labels. The overall architecture of our training framework is illustrated in the lower portion of Figure 2. The following sections describe the key components of our approach: distillation warm-up, policy optimization, and reward calculation.
2.2.1 Distillation for Warm-Start
@@
Due to the limited parameter size of our compression model, its capability for question-focused document summarization is constrained. To provide a strong initial policy for RL and ensure training stability, we employ knowledge distillation from a teacher model to initialize our compressor. Specifically, we first utilize a large-scale language model (DeepSeek-V3) as the teacher to generate summaries of retrieved documents related to the given question. We then evaluate the performance of the downstream LLM ($M$) on the QA task under two conditions: (1) with the teacher-generated summary $\hat{s}$ prepended to the input question q{q} , and (2) with the original question alone. The corresponding performance scores are denoted as $p_{\text{summary}}$ and $p_{\text{original}}$, respectively. By comparing these results, we retain instances where $p_{\text{summary}}>p_{\text{original}}$, indicating that the summary enhances RAG performance. We also retain cases where $p_{\text{original}}=1$ (i.e., the model produces a fully correct answer without the summary) and $p_{\text{summary}}<p_{\text{original}}$; for these, we set the target summary $\hat{s}$ to an empty string. All other instances are discarded. The resulting filtered and modified dataset is denoted as $\mathcal{X}_{f}$, which is used for supervised fine-tuning of the compression model. The fine-tuning objective is defined as:
[…]
where $\pi_{\theta}(q,D)$ denotes the output of the compression model and $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss. This distillation process yields a robust initialization for RL and promotes stability in subsequent training.
2.2.2 End-to-End Training with RL
Following the distillation phase, the compressor possesses a preliminary compression capability. However, as summaries from even the largest teacher models are not guaranteed to be optimal for the downstream task, further end-to-end optimization is necessary. We therefore formulate this optimization as a reinforcement learning problem. In this framework, the compressor functions as a policy that generates a summary from an input question and its corresponding documents. A reward function, which directly reflects performance on the downstream task (e.g., question-answering accuracy), then evaluates the summary’s quality. The objective is to optimize the compressor’s parameters to maximize the expected cumulative reward, thereby directly aligning its outputs with the downstream task’s objectives.
@@ 2.2.2 End-to-End Training with RL
Specifically, we employ Group Relative Policy Optimization (GRPO) (Shao et al., 2024) algorithm. Unlike Proximal Policy Optimization (PPO), which trains a separate critic model, GRPO estimates the baseline directly from a group of rollouts. Given an existing policy, $\pi_{\theta_{\text{old}}}$, and a reference policy, $\pi_{\theta_{\text{ref}}}$, the GRPO objective maximizes the compressor policy $\pi_{\theta}$ using $G$ rollouts $\tau=\{s_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)$, for each input $x\sim\mathcal{D}$:
$\mathcal{J}(\theta)$ $={\mathbb{E}}_{x\sim\mathcal{D},\ \{s_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}$ (2)
@@
$\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Bigg(\frac{\pi_{\theta}(s_{i}|x)}{\pi_{\theta_{\text{old}}}(s_{i}|x)}A_{i},\ \text{clip}\Bigg(\frac{\pi_{\theta}(s_{i}|x)}{\pi_{\theta_{\text{old}}}(s_{i}|x)},1-\epsilon,1+\epsilon\Bigg)A_{i}\Bigg)-\beta\mathbb{D}_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\theta_{\text{ref}}}\big)\Bigg],$
@@
2.2.3 Reward Calculation
  • Generating End-Task Output. It is important to note that the reward is not computed directly from the compressor’s output summary. Instead, the summary $s$ generated by the compressor is prepended to the original input question $q$, and this combined input is fed into the LLM $M\colon(s,q)\mapsto\hat{y}$ to produce a predicted answer $\hat{y}$. The reward is then calculated by comparing $\hat{y}$ to the gold answer $y$. Throughout the training process, the parameters of $M$ remain fixed and are not updated.
@@ 2.2.3 Reward Calculation
    • EM Reward ($r_{\text{EM}}$). We employ EM as the main reward function, which is a widely adopted metric for evaluating the accuracy of QA tasks. The EM reward is binary: it yields a value of 1 if the generated answer perfectly matches the ground truth, and 0 otherwise.
      $r_{\text{EM}}=\begin{cases}1&\text{if }y=\hat{y},\\ 0&\text{otherwise}.\end{cases}$ (3)
    • F1 Reward ($r_{\text{F1}}$). Since exact matches occur infrequently in practice, relying solely on EM rewards leads to sparse reward signals. Furthermore, the EM metric fails to distinguish between partially correct answers, as all non-exact matches receive zero reward. Therefore, we introduce F1 reward, which provides a finer-grained evaluation by measuring the degree of partial match.
      $r_{\text{F1}}=\frac{2\times I_{N}}{P_{N}+R_{N}},$ (4)
      where $P_{N}$ denotes the number of tokens in the predicted answer, $R_{N}$ denotes the number of tokens in the gold answer, and $I_{N}$ is the number of intersecting tokens between the two answers.
    The final reward function combines these reward signals through weighted summation:
@@
    where $\alpha\in(0,1]$ is a hyperparameter that controls the relative contribution of the F1 reward.
2.2.4 Training Template
Figure 5 displays the prompt employed to train the compressor model for generating a summary of the retrieved documents, conditioned on the given question. This prompt is notably concise. For end-task answer generation, the prompt provided to the LLM $M$ is illustrated in Figure 6, which incorporates few-shot in-context examples, the (generated summary of) retrieved documents, and the question.
@@ 3.1 Experimental Settings
  • Datasets and Evaluation Metrics. We evaluate our method on four benchmark datasets: two single-hop question-answering datasets, Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), as well as two multi-hop question-answering datasets, HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). Results are reported on the test sets of Natural Questions and TriviaQA, as well as the development sets of HotpotQA and 2WikiMultihopQA. Following RECOMP (Xu et al., 2024), the performance is measured using Exact Match and token-level F1 scores, while efficiency is assessed by the number of tokens provided in the context.
@@
  • Compression Model ($\pi_{\theta}$). We trained our compression model using Qwen2.5-1.5B-Instruct to generate summaries of the retrieved documents. To evaluate the effect of using different models as compressors, we also trained compressors using Llama3.2-1B-Instruct and Llama3.2-3B-Instruct (Section 3.3).
  • Large Language Model ($M$). We use Qwen2.5-14B-Instruct as the primary LLM model $M$ to generate predicted answers which are used to guide the training of the compressor. To evaluate the generalization ability of our method, we also transfer to another LLM model, LLama3.1-8B-Instruct (Table 2).
@@
  • Retrieval Corpus and Retrievers. Following previous studies (Xu et al., 2024), we use the Wikipedia corpus from December 20, 2018, as the retrieval source for all four datasets. The articles are segmented into non-overlapping 100-word documents. To ensure that our method is not dependent on a specific retriever, we experiment with several mainstream retrievers. Specifically, we use DPR (Karpukhin et al., 2020) for NQ, a hybrid of DPR and BM25 (Robertson et al., 1995) for TriviaQA, and the Contriever model (Izacard et al., 2021) trained on the MS MARCO dataset (Nguyen et al., 2016) for HotpotQA and 2WikiMultihopQA.
@@
  • Baselines. To evaluate the effectiveness of our method, we compared it against various baselines. First, we evaluated the uncompressed approach—retaining the original in-context RALM setup—by prepending the top 1, 3, 5, and 10 retrieved documents to the prompt. We also tested alternative compression methods, including the traditional BM25 algorithm (which ranks sentences by their similarity to the input), off-the-shelf Qwen2.5-1.5B-Instruct model (with comparable parameter size to our approach), and DeepSeek-V3 model (671B parameters, far exceeding our compressor’s capacity). Furthermore, we included state-of-the-art context compression methods for RAG, RECOMP (Xu et al., 2024) and NoiseFilter-IB (Zhu et al., 2024). For RECOMP, we evaluated both its abstractive and extractive variants. To ensure a fair comparison, RECOMP, NoiseFilter-IB, and our approach were all trained using the same model.
@@
  • Implementation Details. For the distillation warm-up phase, we perform full-parameter supervised fine-tuning on the off-the-shell language model for two epochs using LLaMA-Factory  [^1] https://github.com/hiyouga/LLaMA-Factory. This warmed-up model then serves as the initializer for the subsequent reinforcement learning phase. We adopt the Verl framework  [^2] https://github.com/volcengine/verl for RL training. The initialized compression model is trained for two epochs on each dataset. Training is conducted on eight NVIDIA H20 GPUs using full parameter GRPO optimization, with a learning rate of 1e-5, a batch size of 256, five rollouts per sample, and a KL loss coefficient of 0.001. The downstream LLM ($M$) used for reward generation is served using the vLLM inference engine during RL training.
@@
Table 1: Open-domain QA results using Qwen2.5-14B-Instruct as the downstream LLM ($M$). The reported token counts represent the length of in-context documents, excluding few-shot examples. RECOMP, NoiseFilter-IB and our method CORE are all trained using Qwen2.5-1.5B-Instruct.